How many calories of heat were added to 246.9 g of water to raise its temperature from 25oC to 55oC?
1
Expert's answer
2020-02-26T05:01:42-0500

We solve by the formula:

Q=c*m*ΔT=4.184 J/g °C*246.9 g*(55°C-25°C)=30 990.888 J=7.40 kcal

Q - absorbed heat

c- specific heat of water

ΔT - temperature change

The basis of the translation of quantities: 1 kcal = 4186.8 J